i can't remember how to solve these help please
7 - 5x +8 + 2x + 4x - 3 = 2 + 3*5
Thanks

Answer by checkley71(8403)

first you need to combine like terms
-5x+2x+4x=-7-8+3+2+3*5
x=-10+35
x=25 answer.